# Compactifications of phylogenetic systems and electrical networks

### v3

@@ 2.2. Phylogenetic Split Systems
A split of $[n]$ is a bipartition $A|B$ of $[n]$ and a split system is any collection of splits of $[n]$. A graph with some of its nodes labeled by $[n]$ displays a split $A|B$ if there is a set of edges whose removal increases the number of connected components by one, and the two new components include respectively the nodes labeled by $A$ and $B$. A minimal display of the split is a displaying set of edges that do not contain a proper subset displaying that split. A split network is a representation of a split system as a graph, where each split is minimally displayed by a set of parallel edges of the same length. A circular split system can be drawn as a split network with the $n$ boundary nodes on the boundary of a disk, and with non-crossing sets of parallel edges. The following space appears as a single chamber in a larger space of networks, those allowing any ordering of their boundaries, in [11]:
@@
Let $\mathbf{\Psi}_{n}$ be the space of weighted circular split systems on $n$ boundary nodes, using the clockwise cyclic ordering of $[n]$.
The space $\mathbf{\Psi}_{n}$ is stratified with cells corresponding to split systems with no edge weighting. The total number of cells is $|\mathbf{\Psi}_{n}|=2^{n\choose 2}$. A cell $x$ is contained in another cell $y$ when the system corresponding to $x$ can be found by removing splits of $y$. In fact, as a CW-complex $\mathbf{\Psi}_{n}$ is equivalent to a positive orthant of ($\mathbb{R}^{n\choose 2}$. The right side of Figure 4 shows the cells of $\mathbf{\Psi}_{4}$, each column corresponds to cells of a fixed dimension, with $f$-vector $(1,6,15,20,15,6,1)$, the 6th row of Pascal’s triangle.
Remark.
A circular split system can also be represented by a polygonal diagram [11]: the $n$ terminals label the edges of an $n$-gon in some cyclic order, and each split is drawn as a diagonal. Figure 3 displays a polygonal diagram in (c) of the circular split network drawing in (d).
To any circular split system $n$, we can associate a dissimilarity matrix, an $n\times n$ real, symmetric, nonnegative matrix, where the $(ij)$-entry is the sum of the weighted splits between nodes $i$ and $j$ in the network. In our case, the dissimilarity matrix is the resistance matrix $W$. It is well-known [38] that there is a bijection between weighted circular split systems and dissimilarity matrices satisfying the Kalmanson condition: there exists a cyclic order of the boundary such that for any four nodes $i,j,k,l$ listed in that cyclic order,
$\max(W_{ij}\ +\ W_{kl},\ W_{jk}\ +\ W_{il})\ \leq\ W_{ik}\ +\ W_{jl}\,.$
Kalmanson proved that the metrics obeying this condition yield solutions to the Traveling Salesman Problem in polynomial time [27] . Thus, for a fixed circular labeling, the space of Kalmanson dissimilarity matrices of circular split systems on this labeling can be identified with $\mathbf{\Psi}_{n}$.
@@ 3.1. Graphical Map $\mathbf{\xi}$
We define a key function which takes any equivalence class of circular electrical networks as input and returns a circular planar split system. The input circular networks (and their response matrices) must be with respect to a given cyclic ordering, typically the clockwise order. While all our maps are defined on weighted networks (edges weighted with conductance and output weighted split systems), we often do abuse notation and apply the maps to unweighted input and output to see their purely combinatorial function.
@@
Here, each edge of $K(N)$ is reinterpreted as a split whose weight is equal to the conductance of the edge of $K(N).$ Since any response matrix $M$ is seen as a Kron reduction, it is clear that $\mathbf{\xi}$ is injective and surjective, and a homeomorphism to the space of circular split systems with clockwise ordering. Indeed on the response matrices, it is just the identity map together with alternate interpretation of the effective conductance between node $i$ and $j$, or upper-triangular (off-diagonal) entry $M_{i,j}$. [^2] For instance, if $n=8$, $M_{3,8}$ is the weight of the split $\{4,5,6,7,8\}|\{1,2,3\}$. Note that this immediately gives rise to a Kalmanson metric $\mathbf{d}_{\mathbf{\xi}}$ on $[n]$, where the distance $\mathbf{d}_{\mathbf{\xi}}(i,j)$ between $i$ and $j$ is the sum of the splits that separate $i$ from $j.$
@@
Interestingly, while this metric is the effective resistance $W(N^{*})$ for the dual electric network $N^{*}$ for circular planar networks $N$, no measurable interpretation of the metric $\mathbf{d}_{\mathbf{\xi}}$ exists in general.
@@
Thus the action of $\mathbf{\xi}$ on the space of all electrical networks (all response matrices, using clockwise order) is an injective map which is surjective onto the space $\mathbf{\Psi}_{n}$ of circular planar split systems in clockwise order. However the action of $\mathbf{\xi}$ on the cell structure of $\mathbf{\Omega}_{n}$ is an embedding into the cell structure of $\mathbf{\Psi}_{n}.$ Figure 4 shows the combinatorial action of $\mathbf{\xi}$ on the cells of the space $\mathbf{\Omega}_{4}.$ The map $\mathbf{\xi}$ takes all four networks (highlighted region at the top) to the single (highlighted region at the bottom) top-dimensional cell of $\mathbf{\Psi}_{4}$. The split systems not inside the dashed lines are not in the image of $\mathbf{\xi}.$
@@ 3.2. Kalmanson Map $\mathbf{\rho}$
Results in [13] and [14] introduced the correspondence between circular planar electrical networks and circular split systems. The latter paper proved the case for level-1 networks and the former for all connected circular planar electrical networks. Here, we extend that to the case of multiple connected components:
@@
If a symmetric matrix $M$ is a response matrix $M=M(N)$ for a circular planar electrical network $N$, then its resistance matrix $W(N)$ obeys the Kalmanson condition.
Proof.
The case for connected networks can be proven using either Kirchhoff and Ohm’s laws [13] or Kenyon and Wilson’s equation [29, Theorem 2.9]. The latter equation also applies to the disconnected case, where the resistance matrix will have zeros, which correspond to nodes with no split between them. ∎
Theorem 8 implies that for any circular planar electrical network, there is a corresponding circular split system. The construction of that system is via the Buneman algorithm [38] or the Neighbor-Net algorithm [5]. The former requires precise resistances, but the latter can work with approximates. For accurate data, they both give the same result, which we define as follows:
@@
The Kalmanson split system $\mathbf{\rho}(N)$ is the circular split system corresponding (injectively) to the (equivalence class of) circular planar electrical network $N$ [^3] The Kalmanson map we call $\mathbf{\rho}$ here is just the extension of the map $R_{w}$ defined in [13], from connected planar networks to all planar networks.
The Kalmanson map $\mathbf{\rho}$ can be extended to all (non-planar) electrical networks by letting even non-Kalmanson resistance metrics be subjected to the best-fit approximation of Neighbor-net. In that extension, the map will no longer be injective. Note also that the algorithm will find a cyclic order for which $W(N)$ is Kalmanson if possible. This output will only have the original cyclic order if the original is circular planar in that order.
@@ 3.3. Induced Map $\mathbf{\sigma}$
For the case of a connected circular planar electrical network as input, we define the induced split system. [^4] For electrical networks with multiple connected components, we must wait until we define the proper codomain — the compactification of the space of split systems — in Section 4. For a circular planar electrical network $N$, a grove is a spanning forest of the graph of $N$ whose component trees each include some of the nodes labeled by $[n].$ A $k$-grove is a grove with $k$ trees. The weight of a grove is the product of the weights of all the edges in that grove. A 2-grove of $N$ respects a split that is minimally displayed by $N$ if the two trees of the grove span the two components of the displayed split.
@@
When restricted to circular planar electrical networks $N$, the induced split system coincides with the Kalmanson split system. That is, $\mathbf{\rho}(N)=\mathbf{\sigma}(N)$.
@@
We prove this for the case of a connected network $N$. For a network with more than one connected component, the proof applies to each separate component, via block-decomposition of the matrices. Consider a connected circular planar electrical network $N$ which displays the splits $\mathbf{\sigma}(N)$. Each split corresponds to a collection of the 2-groves that Kenyon and Wilson use in their formula for the resistance $R_{ij}$: the 2-groves that are possible after deleting any set of edges that display the split [29, Proposition 2.7]. Note that this correspondence partitions the 2-groves. Thus the sum of weights of the splits between two nodes is the same as the sum of the weights of the 2-groves. ∎
@@ 3.4. Matchings and Duals
Although the graphical map at first seems merely a visual coincidence, the next theorem shows that it actually extends our earlier maps to all circular networks, planar and nonplanar. Thus it exhibits all circular electrical networks (up to equivalence) in a one-to-one correspondence with all circular planar split systems, with the planar circular electrical networks embedded as a special subset.
@@
Again we prove this for the case of a connected network $N$. For a network $N$ with more than one connected component, the response matrix will decompose into blocks on the diagonal for each component. The resistance between components would be infinite, so mapping must be done component-wise, where the function $\mathbf{\rho}$ (finding the resistance $W$ and the corresponding splits) can be applied for each component via its block matrix. [^7] For a network with more than one connected component, the split system we obtain will have multiple components, lying in the compactification of $\mathbf{\Psi}_{n}$; see Section 4.
Proof of Theorem 13.
@@
3.5. Summary of Maps
Figure 7. Domains and ranges for the maps in this paper.
This construction raises new questions. First, although here we restrict our study to planar networks with a given clockwise order of nodes, the map $\mathbf{\xi}$ can operate on any network, as long as the nodes are given a clockwise order first. The results of that mapping for non-planar examples are interesting, but not studied here. Moreover, the general relationship between the weights of splits to their original conductances remains unknown.
Secondly the map $\mathbf{\xi}$ carries the structure of duality on electrical networks to a new duality on their image in the circular split systems. We note that the map $\mathbf{\xi}$ is reminiscent of the $T$-duality map on decorated permutations [35]. Moreover, $\mathbf{\xi}$ carries an equivalence relationship on circular split systems — under twisting as defined by Devadoss and Petti [11] — to a new combinatorial equivalence on response matrices which rewires the (possibly nonplanar) electrical networks. What is the physical meaning of this equivalence? Indeed, $\mathbf{\xi}$ shows the poset and CW cell structures on circular split systems as a coarsening of those on circular planar networks. It also carries the cell structure of split systems to a new cell structure on all circular electrical networks. This new combinatorial structure is unexplored.
@@ 4.1. Cactii
A cactus electrical network is a generalized circular planar network where boundary nodes are allowed to be identified. In particular, the space of cactus networks is the compactification of the space of circular planar networks: the conductance between two boundary nodes is allowed to become $\infty$ (thereby “short-circuiting” them). Figure 8(a) shows an example along with its set of splits, a construction described in [23, 34]. The cactus network $N$ has an associated strand diagram $P(N)$ in part (b), which now can have several boundary nodes occupying the same shaded region. Again we use the medial graph of $N$. However, to draw the strands, place all the nodes in clockwise order on a single circular boundary. Include semi-circular strands around any disconnected boundary node, and require that any set of boundary nodes all occupying the same shaded region are a set identified in the network; see part (c). Just as in Definition 12, the strand diagram allows the construction of the planar dual $N^{*}$ for a cactus network seen in Figure 8(d).
Figure 8. Cactus network $N$, strand diagrams, and its dual $N^{*}$.
Two cactus networks are equivalent if they have the same sets of response matrices, one for each bulb. Lam [34] shows that the (unweighted) electrical equivalence classes of circular planar electrical networks, including those formed by identifying terminal points, correspond bijectively to the perfect matchings on $[2n]$. To find the perfect matching guaranteed by this bijection, the equivalence class must be represented by a reduced $N$. In that case, the strands of the strand diagram will obey the requirement that any two of them cross each other at most once.
@@
For a given cyclic order on $[n]$, the space of equivalence classes of cactus networks is the compactification ${\mathbf{\Omega}}_{n}$ of the space of circular electrical networks $\mathbf{\Omega}_{n}$. [^8] In [34], the space ${\mathbf{\Omega}}_{n}$ is denoted as $E_{n}$.
@@
Figure 9. The graphical map $\mathbf{\xi}$ from cactus electrical networks ${\mathbf{\Omega}}_{3}$ to cactus split systems $\mathbf{\Psi}_{3}$

### v4

@@ 2.2. Phylogenetic Split Systems
A split of $[n]$ is a bipartition $A|B$ of $[n]$ and a split system is any collection of splits of $[n]$. A graph with some of its nodes labeled by $[n]$ displays a split $A|B$ if there is a set of edges whose removal increases the number of connected components by one, and the two new components include respectively the nodes labeled by $A$ and $B$. A minimal display of the split is a displaying set of edges that do not contain a proper subset displaying that split. A split network is a representation of a split system as a graph, where each split is minimally displayed by a set of parallel edges of the same length. A circular split system can be drawn as a split network with the $n$ boundary nodes on the boundary of a disk, and with non-crossing sets of parallel edges. A circular split system can also be represented by a polygonal diagram [12]: the $n$ terminals label the edges of an $n$-gon in some cyclic order, and each split is drawn as a diagonal. Figure 1 has an unweighted polygonal diagram in (c) which shows the split system on $[4]$ with all four trivial splits (these have one singleton set) and the split $\{1,4\}|\{2,3\}.$ Figure 4 displays a weighted polygonal diagram in (c) of the circular split network drawing in (d). Following [12] we weight the splits by assigning each split a non-negative real value independently. The weight of zero is equivalent to removing that split entirely.
@@
Let $\mathbf{\Psi}_{n}$ be the space of weighted circular split systems on $n$ boundary nodes, using the clockwise cyclic ordering of $[n]$. Since their are (${n\choose 2}$ splits possible, the resulting space is isomorphic to the nonnegative orthant of $\mathbb{R}^{n(n-1)/2}.$
The space $\mathbf{\Psi}_{n}$ is stratified with cells corresponding to split systems with no edge weighting. The total number of cells is $|\mathbf{\Psi}_{n}|=2^{n\choose 2}$. A cell $x$ is contained in another cell $y$ when the system corresponding to $x$ can be found by removing splits of $y$. In fact, as a CW-complex $\mathbf{\Psi}_{n}$ is equivalent to a nonnegative orthant of ($\mathbb{R}^{n\choose 2},$ with cells found as the regions of the coordinate hyperplanes, axes, etc. The right side of Figure 5 shows the cells of $\mathbf{\Psi}_{4}$, each column corresponds to cells of a fixed dimension, with $f$-vector $(1,6,15,20,15,6,1)$, the 6th row of Pascal’s triangle.
Next we review, that to any circular split system $n$, we can associate a dissimilarity matrix, an $n\times n$ real, symmetric, nonnegative matrix, where the $(ij)$-entry is the sum of the weighted splits between nodes $i$ and $j$ in the network. In our case, the dissimilarity matrix is the resistance matrix $W$. It is well-known [39] that there is a bijection between weighted circular split systems and dissimilarity matrices satisfying the Kalmanson condition: there exists a cyclic order of the boundary such that for any four nodes $i,j,k,l$ listed in that cyclic order,
(2.2) $\max(W_{ij}\ +\ W_{kl},\ W_{jk}\ +\ W_{il})\ \leq\ W_{ik}\ +\ W_{jl}\,.$
Thus, for a fixed circular labeling, the space of Kalmanson dissimilarity matrices of circular split systems on this labeling can be identified with $\mathbf{\Psi}_{n}$. Kalmanson proved that the metrics obeying this condition yield solutions to the Traveling Salesman Problem in polynomial time [28] .
@@ 3. Maps between Classical Spaces
In this section we explain functions that take input the connected circular (planar) electrical networks. The functions will be extended to multiple connected components and cactus versions of the electrical networks in Section 4, after we discuss duals on cactus networks and define cactus split systems.
@@ 3.1. Graphical Map $\mathbf{\xi}$
We start by defining a key function which takes any equivalence class of circular electrical networks (planar or non-planar) as input and returns a circular planar split system. The input circular networks (and their response matrices) must be planar with respect to a given cyclic ordering, typically the clockwise order. While all our maps are defined on weighted networks (edges weighted with conductance), and the outputs are weighted split systems, we often do abuse notation and apply the maps to unweighted input and output to see their purely combinatorial function.
@@
Here, each edge of $K(N)$ is reinterpreted as a split whose weight is equal to the conductance of the edge of $K(N).$ Since any response matrix $M$ is seen as a Kron reduction, it is clear that $\mathbf{\xi}$ is injective and surjective, and a homeomorphism to the space of circular split systems with clockwise ordering. Indeed, seen as a function on the response matrices, it is just the identity map (on the upper triangular submatrix, which determines the rest of the response matrix), together with alternate interpretation of the effective conductance between node $i$ and $j$, the entry $M_{i,j}$. [^2] For instance, if $n=8$, $M_{3,8}$ is the weight of the split $\{4,5,6,7,8\}|\{1,2,3\}$. Note that this immediately gives rise to a Kalmanson metric $\mathbf{d}_{\mathbf{\xi}}$ on $[n]$, where the distance $\mathbf{d}_{\mathbf{\xi}}(i,j)$ between $i$ and $j$ is the sum of the splits that separate $i$ from $j.$
@@
Interestingly, while we will see that this metric is the effective resistance $W(N^{*})$ for the dual electric network $N^{*}$ for circular planar networks $N$, no meaningful interpretation of the metric $\mathbf{d}_{\mathbf{\xi}}$ is known for general, non-planar inputs.
@@
Thus the action of $\mathbf{\xi}$ on the space of all electrical networks (all response matrices, using clockwise order) is an injective map which is surjective onto the space $\mathbf{\Psi}_{n}$ of circular planar split systems in clockwise order. However the action of $\mathbf{\xi}$ on the cell structure of $\mathbf{\Omega}_{n}$ is an embedding into the cell structure of $\mathbf{\Psi}_{n}.$ Figure 5 shows the combinatorial action of $\mathbf{\xi}$ on the cells of the space $\mathbf{\Omega}_{4}.$ The map $\mathbf{\xi}$ takes all four networks (highlighted region at the top) to the single (highlighted region at the bottom) top-dimensional cell of $\mathbf{\Psi}_{4}$. Even when restricted to that cell, the action of $\mathbf{\xi}$ is not onto: only certain weighted split systems with all 6 nonzero splits arise from circular planar electrical networks. Let the values of splits be $a,b,c,d$ for the trivial splits separating 1,2,3,4 in that order, and the values be $e,f$ for the interior splits. Then the range of $\mathbf{\xi}$ inside the set of split systems (with all 6 splits nonzero) are those satisfying $ac\geq ef$ and $bd\geq ef.$ (This follows from the non-negativity of circular minors in the response matrix.) All the other split systems inside the dashed lines are cells for which $\mathbf{\xi}$ is onto. The split systems not inside the dashed lines are not in the image of $\mathbf{\xi}$ at all.
@@ 3.2. Kalmanson Map $\mathbf{\rho}$
Results in [14] and [15] introduced the correspondence between circular planar electrical networks and circular split systems. The latter paper proved the case for level-1 networks and the former for all connected circular planar electrical networks. Here we review those results, and in Section 4 we extend it that to the cases of cactus networks and multiple connected components. The following is proven in [14]:
@@
If a symmetric matrix $M$ is a response matrix $M=M(N)$ for a connected circular planar electrical network $N$, then its resistance matrix $W(N)$ obeys the Kalmanson condition.
Theorem 8 implies that for any connected circular planar electrical network, there is a corresponding circular split system. The construction of that system is via the Buneman algorithm [39] or the Neighbor-Net algorithm [6]. The former requires precise resistances, but the latter can work with approximates. For accurate data, they both give the same result, which we define as follows:
@@
The Kalmanson split system $\mathbf{\rho}(N)$ is the circular split system corresponding (injectively) to the Kalmanson resistance metric $W(N)$ of the (equivalence class of the) circular planar electrical network $N$ [^3] The Kalmanson map we call $\mathbf{\rho}$ here is the map $R_{w}$ defined in [14], but we will extend it from connected planar networks to all cactus networks.
The Kalmanson map $\mathbf{\rho}$ can be extended to non-planar electrical networks by letting even non-Kalmanson resistance metrics be subjected to the best-fit approximation of Neighbor-net. In that extension, the map will no longer be injective. Note also that the algorithm will find a cyclic order for which $W(N)$ is Kalmanson, if possible. This output split system will only have the original cyclic order if the original is circular planar in that order.
@@ 3.3. Induced Map $\mathbf{\sigma}$
For the case of a connected circular planar electrical network as input, we define the induced split system. For a circular planar electrical network $N$, a grove is a spanning forest of the graph of $N$ whose component trees each include some of the nodes labeled by $[n].$ A $k$-grove is a grove with $k$ trees. The weight of a grove is the product of the weights of all the edges in that grove. A 2-grove of $N$ respects a split that is minimally displayed by $N$ if the two trees of the grove span the two components of the displayed split.
@@
For connected circular planar electrical networks $N$, the induced split system coincides with the Kalmanson split system. That is, $\mathbf{\rho}(N)=\mathbf{\sigma}(N)$.
@@
Consider a connected circular planar electrical network $N$ which displays the splits $\mathbf{\sigma}(N)$. Each split corresponds to a collection of the 2-groves that Kenyon and Wilson use in their formula for the resistance $R_{ij}$: the 2-groves that are possible after deleting any set of edges that display the split [30, Proposition 2.7]. Note that this correspondence partitions the 2-groves. Thus the sum of weights of the splits between two nodes is the same as the sum of the weights of the 2-groves. ∎
@@ 3.4. Matchings and Duals
Although the graphical map at first seems merely a visual coincidence (Edges in the Kron reduction look like splits in a circular split system!), the next theorem shows that it actually extends our earlier maps to all circular networks, planar and non-planar. Thus it exhibits all circular electrical networks (up to equivalence) in a one-to-one correspondence with all circular planar split systems, with the planar circular electrical networks embedded as a special subset.
@@
Again we prove this for the case of a connected network $N$.  [^6] For a network with more than one connected component, the split system we obtain will have multiple components, lying in the compactification of $\mathbf{\Psi}_{n}$; see Section 4.
Proof.
@@ 4.1. Cactii
A cactus electrical network is a generalized circular planar network where boundary nodes are allowed to be identified. In particular, the space of cactus networks is the compactification of the space of circular planar networks. In a cactus network the conductance between two boundary nodes is allowed to become $\infty$ (thereby “short-circuiting” them). This is pictured by pinching together the boundary at those points, creating bulbs of the network connected to each other by nodes that each correspond to an equivalence class of the original terminals. The space of cactus networks was described originally in [34] with more examples in [24]. Figure 8(a) shows an example along with its set of splits. Notice that in our pictures, we include empty bulbs at each multi-identified terminal node. All the original terminal nodes $1,\dots,n$ have labels which are kept in their original clockwise order, with one bulb between each pair of nodes that are identified. This arrangement, including the empty bulbs, is important for seeing the action of our maps to the split systems.
The cactus network $N$ has an associated strand diagram $P(N)$ in part (b), which can have several boundary nodes occupying one shaded region—those which are identified in the network. Again we use the medial graph of $N$. However, to draw the strands, place all the nodes in clockwise order on a single circular boundary. Include semi-circular strands around any disconnected boundary node, and require that any set of boundary nodes all occupying the same shaded region are a set identified in the network; see part (c). Just as in Definition 12, the strand diagram allows the construction of the planar dual $N^{*}$ for a cactus network seen in Figure 8(d).
Figure 8. Cactus network $N$, strand diagrams, and its dual $N^{*}$. The splits displayed by $N$ are shown as shaded cuts in part (a). Notice that in the dual $N^{*}$, the nodes comprising connected components all share the same bulb in $N.$
Each bulb of a cactus network is assigned its own response matrix, with rows and colums indexed by the (identified classes of) terminals on the boundary of that bulb. Two cactus networks are equivalent if they have the same sets of response matrices, one for each bulb. Lam [34] shows that the (unweighted) electrical equivalence classes of circular planar electrical networks, including those formed by identifying terminal points, correspond bijectively to the perfect matchings on $[2n]$. To find the perfect matching guaranteed by this bijection, the equivalence class must be represented by a reduced $N$. In that case, the strands of the strand diagram will obey the requirement that any two of them cross each other at most once.
@@
For a given cyclic order on $[n]$, the space of equivalence classes of cactus networks is the compactification ${\mathbf{\Omega}}_{n}$ of the space of circular electrical networks $\mathbf{\Omega}_{n}$. [^7] As mentioned earlier, in [34], the space ${\mathbf{\Omega}}_{n}$ is denoted as $E_{n}$.
@@
Figure 9. The graphical map $\mathbf{\xi}$ from cactus electrical networks ${\mathbf{\Omega}}_{3}$ to cactus split systems $\mathbf{\Psi}_{3}$. Cells with $k$ edges (or splits) are $k$-dimensional; there are three 2-cells in each picture, one is on the back of the 3-cell.